# A Multimodal Seq2Seq Transformer for Predicting Brain Responses to Naturalistic Stimuli

### v1

@@ Abstract
The Algonauts 2025 Challenge called on the community to develop encoding models that predict whole-brain fMRI responses to naturalistic multimodal movies. In this submission, we propose a sequence-to-sequence Transformer that autoregressively predicts fMRI activity from visual, auditory, and language inputs. Stimulus features were extracted using pretrained models including VideoMAE, HuBERT, Qwen, and BridgeTower. The decoder integrates information from prior brain states, current stimuli, and episode-level summaries via dual cross-attention mechanisms that attend to both perceptual information extracted from the stimulus as well as narrative information provided by high-level summaries of narrative content. One core innovation of our approach is the use of sequences of multimodal context to predict sequences of brain activity, enabling the model to capture long-range temporal structure in both stimuli and neural responses. Another is the combination of a shared encoder with partial subject-specific decoder, which leverages common structure across subjects while accounting for individual variability. Our model achieves strong performance on both in-distribution and out-of-distribution data, demonstrating the effectiveness of temporally-aware, multimodal sequence modeling for brain activity prediction. The code is available at https://github.com/Angelneer926/Algonauts_challenge.
@@ 1 Introduction
To account for the rich, hierarchical structure of naturalistic stimuli, we extracted stimulus features from state-of-the-art pretrained models across multiple modalities: VideoMAE for temporal dynamics of visual motion [8], HuBERT for acoustic features [9], and Qwen for linguistic representations. Additionally, we incorporate sentence-level semantic features derived from BERT [10], using manually sourced high-level summaries of the TV episodes featured in training data. This allows us to provide broader narrative context beyond the moment-to-moment stimulus input [11]. To capture joint visual-linguistic representations, we extract cross-modal features from the stimuli using BridgeTower [12].
@@ 4.1 Feature Extraction
To complement the unimodal features, we utilized BridgeTower, a pretrained cross-modal transformer architecture designed to integrate visual and textual signals. At each time point, the model receives a video frame and its corresponding utterance as input. We extract a 1536-dimensional fused representation from the pooler__\__output of the final layer, which corresponds to the [CLS] token after visual/text fusion. These features provide additional cross-modal alignment that is difficult to obtain from independent visual and language encoders, and serve as a valuable supplement to the modality-specific representations.
Episode Summary.
While the original dataset provided dialogue-level annotations, the episodes contained overarching narrative structures that are essential for interpretation. To compensate for this missing context, we manually sourced episode-level summaries from online resources (e.g., Google searches). Each summary is segmented into sentences and passed through a pretrained BERT model to obtain sentence-level embeddings. We explored two strategies to incorporate episode-level language context into the frame-wise representation learning process. The first is a Gaussian-weighted sum of sentence embeddings based on their temporal distance to each frame. The second is a cross-attention mechanism in the decoder, where the fMRI prediction at each time step attends to the sentence-level embeddings from the episode summary. While the Gaussian-weighted approach imposes a smooth and interpretable temporal bias, cross-attention offers greater flexibility by allowing the model to learn dynamic alignments between frames and narrative content. Although the alignment between summaries and frames was approximate, this additional modality captures high-level thematic information that may guide interpretation, particularly in long-form episodes.
@@ 5 Results
The performance gap between Friends and Movie10 is likely due to the training set containing a larger proportion of Friends data, allowing the model to more effectively learn patterns in that stimulus set. In contrast, Movie10 clips differ substantially in both visual style and narrative structure, posing a greater challenge for generalization. The lower performance on Movie10 is thus expected and reflects the challenge of transferring representations learned from episodic television from a single series to a diverse set of feature films. Another contributing factor may be the absence of episode-level summaries for the movies, which limits the model’s access to high-level semantic context that it could leverage during Friends episodes.
@@
Our model achieved an average Pearson correlation of 0.305 on Friends Season 7 and 0.199 on the OOD movie set. See Figure 6. The strong performance on Season 7 indicates that the model generalizes reliably within the domain it was primarily trained on. Similar to our validation results, performance on the movies were substantially lower, indicating that the model did not transfer as effectively to out-of-distribution stimuli.
@@ 6 Conclusion
In this work, we present a multimodal sequence-to-sequence Transformer model for predicting full-brain fMRI responses to naturalistic audiovisual and linguistic stimuli. By leveraging pretrained encoders across multiple modalities, integrating semantic-level context from episode summaries, and modeling brain activity autoregressively with subject-specific adaptations, our approach captures both the temporal and individual complexity of neural responses. The model demonstrates strong performance on both in-distribution and out-of-distribution data, highlighting the benefits of temporally-aware, cross-modal integration. Moreover, the sliding window training strategy, combined with teacher forcing, enables efficient learning despite limited data. Future directions include incorporating more precise alignment between high-level semantic context and neural dynamics, scaling the model to broader subject cohorts, and exploring zero-shot generalization across unseen individuals and stimuli. Our results support the promise of generative, multimodal models as a powerful framework for bridging perception and brain activity in real-world settings.

### v2

@@ Abstract
The Algonauts 2025 Challenge called on the community to develop encoding models that predict whole-brain fMRI responses to naturalistic multimodal movies. In this submission, we propose a sequence-to-sequence Transformer that autoregressively predicts fMRI activity from visual, auditory, and language inputs. Stimulus features were extracted using pretrained models including VideoMAE, HuBERT, Qwen, and BridgeTower. The decoder integrates information from prior brain states and current stimuli via dual cross-attention mechanisms that attend to both perceptual information extracted from the stimulus as well as narrative information provided by high-level summaries of the content. One core innovation of our approach is the use of sequences of multimodal context to predict sequences of brain activity, enabling the model to capture long-range temporal structure in both stimuli and neural responses. Another is the combination of a shared encoder with partial subject-specific decoder, which leverages common representational structure across subjects while accounting for individual variability. Our model achieves strong performance on both in-distribution and out-of-distribution data, demonstrating the effectiveness of temporally-aware, multimodal sequence modeling for brain activity prediction. The code is available at https://github.com/Angelneer926/Algonauts_challenge.
@@ 1 Introduction
To account for the rich, hierarchical structure of naturalistic stimuli, we extracted stimulus features from state-of-the-art pretrained models across multiple modalities: VideoMAE for temporal dynamics of visual motion [8], HuBERT for acoustic features [9], and Qwen for linguistic representations. Additionally, we incorporated sentence-level semantic features derived from BERT [10], using high-level summaries of the TV episodes and movies that constitute the training and testing data. This allows us to provide broader narrative context beyond the moment-to-moment stimulus input [11]. To capture joint visual-linguistic representations, we extracted cross-modal features from the stimuli using BridgeTower [12].
@@ 4.1 Feature Extraction
To complement the unimodal features, we utilized BridgeTower, a pretrained cross-modal transformer architecture designed to integrate visual and textual signals. At each time point, the model receives a video frame and its corresponding utterance as input. We extract a 1536-dimensional fused representation from the pooler_\__output of the final layer, which corresponds to the [CLS] token after visual/text fusion. These features provide additional cross-modal alignment that is difficult to obtain from independent visual and language encoders, and serve as a valuable supplement to the modality-specific representations.
Narrative Summaries.
While the original dataset provided dialogue-level annotations, the episodes and movies contained overarching narrative structures that are essential for interpretation. To compensate for this missing context, we manually sourced narrative summaries of each episode and movie.. Each summary is segmented into sentences and passed through a pretrained BERT model to obtain sentence-level embeddings. We explored two strategies to incorporate this additional context into the frame-wise representation learning process. The first is a Gaussian-weighted sum of sentence embeddings based on their temporal distance to each frame. The second is a cross-attention mechanism in the decoder, where the fMRI prediction at each time step attends to the sentence-level embeddings from the narrative summary. While the Gaussian-weighted approach imposes a smooth and interpretable temporal bias, cross-attention offers greater flexibility by allowing the model to learn dynamic alignments between frames and narrative content. Although the alignment between summaries and frames was approximate, this additional modality captures high-level thematic information that may guide interpretation, particularly in long-form narratives.
@@ 5 Results
The performance gap between Friends and Movie10 is likely due to the training set containing a larger proportion of Friends data, allowing the model to more effectively learn patterns in that stimulus set. In contrast, Movie10 clips differ substantially in both visual style and narrative structure, posing a greater challenge for generalization. The lower performance on Movie10 is thus expected and reflects the challenge of transferring representations learned from episodic television from a single series to a diverse set of feature films.
@@
Our model achieved an average Pearson correlation of 0.305 on Friends Season 7 and 0.199 on the OOD movie set. See Figure 6. The strong performance on Season 7 indicates that the model generalizes reliably within the domain it was primarily trained on. Similar to our validation results, performance on the movies was lower, indicating that the model did not transfer as effectively to out-of-distribution stimuli.
@@ 6 Conclusion
In this work, we present a multimodal sequence-to-sequence Transformer model for predicting full-brain fMRI responses to naturalistic audiovisual and linguistic stimuli. By leveraging pretrained encoders across multiple modalities, integrating semantic-level context from narrative summaries, and modeling brain activity autoregressively with subject-specific adaptations, our approach captures both the temporal and individual complexity of neural responses. The model demonstrates strong performance on both in-distribution and out-of-distribution data, highlighting the benefits of temporally-aware, cross-modal integration. Moreover, the sliding window training strategy, combined with teacher forcing, enables efficient learning despite limited data. Future directions include incorporating more precise alignment between high-level semantic context and neural dynamics, scaling the model to broader subject cohorts, and exploring zero-shot generalization across unseen individuals and stimuli. Our results support the promise of generative, multimodal models as a powerful framework for bridging perception and brain activity in real-world settings.